SnowyBreaks:

Is it bad to lose 3.8 pounds in less than 2 days?

AsianPanda08:

Nope. that's actually good. unless you've got a condition or ur forcing urself to.

SnowyBreaks:

@asianpanda08 wrote:
Nope. that's actually good. unless you've got a condition or ur forcing urself to.
Mainly high metabolism, but refusing to eat aswell

AsianPanda08:

oh then yea, dats not good

Nina001:

that's not good

Nina001:

come with me

Nina001:

we can be big backs together

Nina001:

but its bad specially if you are skinny

AsianPanda08:

ME TOO

Nina001:

YUHHH

Nina001:

actuallyyy

Nina001:

I think me loosing 3.8 pounds is totally ok...

Nina001:

my butt is js gonna go flat...

Nina001:

smhhhh

AsianPanda08:

i needa lose like 10 pounds

Nina001:

girll

Nina001:

you mad skinny

Nina001:

haven't you seen mee

AsianPanda08:

i'm not :sob:

AsianPanda08:

Yes i've seen u

AsianPanda08:

u showed me

Nina001:

yupp

Nina001:

rate me 1-10

AsianPanda08:

100/10

Nina001:

awhhhh

Nina001:

but you are a 200/10 plus extra credit 240/10

AsianPanda08:

ok then ur 300/10 + 300 = 600/10

Nina001:

nopeeee

AsianPanda08:

gurl bsfr ur actually rlly pretty

Nina001:

you aree 600/10 +800= 1400/10

Nina001:

@asianpanda08 wrote:
gurl bsfr ur actually rlly pretty
CAPPP I'm so uglyyyy

AsianPanda08:

NUH UH

Nina001:

totally ugly and insecuree

AsianPanda08:

WAIT Y R WE CHATTING ON DIS

Nina001:

LOL GETTING LEVEL UP I THINK.... or just because we randomly started talking on the subject

AsianPanda08:

true

Nina001:

yo

Nina001:

me andd my friend were getting our levels up

Nina001:

xshawtyx

Nina001:

and me got to 67 and 27

AsianPanda08:

ooo

AsianPanda08:

on wat

Ferrari:

I have a extremely high metabolism (125 at 5'11). For me, losing that much in 2 days is awful news. I am trying to gain weight lol. If you weigh a bit more than average, losing the weight is a good thing. If you are trying to lose the weight, congrats!

SnowyBreaks:

I weigh 113.